find the number of years it would take for 748 to increase by 72 percent if invested at interest rate 7.9 percent per annum, compounded semi-annually


1
Expert's answer
2021-06-29T12:48:09-0400

Given,

Principal amount = 748

Increased amount = 72 % of 748

So, total amount = "\frac{172\times 748 }{100} =1286.56"

Rate of interest = 7.9% per annum compounded semi annually

So, rate of interest for the semi annual "= \frac{7.9}{2} = 3.45" %


"A= P(1+R\%)^n"

"\Rightarrow 1286.56 = 748(1+\frac{3.45}{100})^n"

"\Rightarrow 1.72=1.0345^n"

Taking log of both side,

"\Rightarrow n= \frac{\log(1.72)}{\log(1.0345)}"

"\Rightarrow n = 15.98"

"\Rightarrow n \simeq 16"

Hence, total number of year will be 8 years.